Group Selection Page
HELP
Controls Group
Data Analysis Group
ITDM
WP76
DET
Vacuum
EEE
WP78
WP74
FXE
SCS
SPB
MID
HED
SQS
SXP
Sample Environment
Operation
PSPO
XO
TS
Migrated2Zulip
|
Technical integration and coordination
SASE1 Status and SASE3 Readiness
|
Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther |
Not logged in |
 |
|
Message ID: 291
Entry time: 12 Jan 2022, 10:06
|
Author: |
Frederik Wolff-Fabris |
Location: |
General |
Severity: |
Information |
Subject: |
Notes XO Integration Meeting 12.01.2022 |
|
|
General Announcement:
- SEPS interlock tests are scheduled for January 21st: starting early morning from 08:00 to 16:00 across all the SASEs;
In parallel D-3 tests the interlocks in all hutches including SXP. An e-mail to instruments was sent and a reply has to be directed to Zunaira A.
- Disabling SEPS interlock yesterday night was not necessary. Further discussion with A. Galler will follow.
CTRL:
- No planned updates.
DA:
- No updates to report.
EEE:
- Yesterday the SA1 vacuum update took place.
- MPS contacted Nerea J. on the SEPS interlocks issue: Nerea J. to be included in further discussions.
VAC:
- Nothing to report.
XPD:
- Preparing for tunnel closure: finishing Hard X-ray PES tests; works in XTD9 are completed.
- Works in XTD6 finishing EPS switches and installing upgrade M-ports in MCP (crate coming on Thursday and update interlock connection).
DET:
- Nothing to report.
ITDM:
- Remind on the Wednesday maintenance on the dCache and GPFS; Access to XFEL related data from Maxwell compute cluster is not guaranteed during the intervention. Please contact ITDM in case of conflicting
with it.
XRO:
- Not present.
LAS:
- Failed to upgrade the SASE1 front ends during Winter Maintenance (company did not come due to Corona regulations).
- 8D report for the Seeder shutdown issue at the start of the maintenance period: A. Galler informs that by the end of this week it shall be ready and distributed.
- Power supply connection: Max L. asks for XO support to follow up with TS.
UNSYS:
- Apple-X undulators are aligned and currently the intersection is being aligned together with air coils. Vacuum shall also be connected permitting to close tunnel tomorrow.
SPB:
- Nothing to report.
FXE:
- Nothing to report.
MID:
- Thanks to SPB group for helping with providing a flange.
HED:
- Apart from the SEPS all is fine.
SCS:
- Not present.
SQS:
- All fine. |